OpenStudy (anonymous):

Hi!!! I have 2 math questions. Solve for w: w+6=2(w-6) Solve for b: 3(b+1)=4b-1

OpenStudy (anonymous):

answer or know how to do it?

OpenStudy (anonymous):

Know how to do it and answer

OpenStudy (anonymous):

w + 6 = 2w - 12 w = 18 3b - 3 = 4b - 1 b = -2

OpenStudy (anonymous):

Can you explain how you got that? Clearly please??

OpenStudy (anonymous):

first you expand everything

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

like 2(w-6) = 2w - 12

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

And...

OpenStudy (anonymous):

then collect like terms and isolate the variable